OpenStudy (superhelp101):

Is x - 8 a factor of the function f(x) = -2x3 + 17x2 - 64? Explain.

OpenStudy (superhelp101):

Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 64. No. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is zero. Therefore, x - 8 is not a factor of f(x) = -2x3 + 17x2 - 64. Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is not zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 64. No. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is not zero. Therefore, x - 8 is not a factor of f(x) = -2x3 + 17x2 - 64.

OpenStudy (superhelp101):

I know it's between the first option and 3rd option :)

OpenStudy (anonymous):

well i divided it and the remainder is 2112 so the 3rd opinion is correct ;)

OpenStudy (anonymous):

and please let me divide it again to be sure :)

OpenStudy (superhelp101):

oh ok, take your time ;) 

OpenStudy (anonymous):

OH sry! i made a mistake !

OpenStudy (superhelp101):

it's okay.

OpenStudy (anonymous):

now the remainder is zero :)

OpenStudy (superhelp101):

thank you for your help ;)

OpenStudy (anonymous):

you're welcome :)